How much of a 5% solution must be added to 400ml of a 10% salt solution to obtain a 7% solution?
Answer by robertb(5830)

0.05x + 0.10*400 = 0.07(x + 400)
0.05x + 40 = 0.07x + 28
12 = 0.02x
x = 600 ml of the 5% solution.